# Diff-DAgger: Uncertainty Estimation with Diffusion Policy for Robotic Manipulation

### v2

@@ IV-C Task Performance
We evaluate the policy performance of methods for state-based and image-based observations. Diff-DAgger method consistently outperforms the baselines in task completion rate across tasks, assuming the same number of expert demonstrations is given. On average, Diff-DAgger achieves a 12.6% improvement over the second-best method in state-based tasks. For visuomotor tasks, Diff-DAgger on average achieves a 15.0% improvement over the other methods. Notably, Diff-DAgger achieves 100% policy performance with only 52 stacking demonstrations using image data, indicating that it can be used to significantly reduce the number of demonstrations required to train an optimal policy.
@@ IV-D Wall-Clock Time for DAgger
We demonstrate these results on the visuomotor stacking task, where Diff-DAgger reduced training time by a factor of 5.1, threshold setting time by a factor of 24.5, and inference time by a factor of 5.3 compared to Ensemble-DAgger, as shown in Fig 2. The overall reduction in average policy update time is by a factor of 5.4, reducing the total time to just 0.37 hours. Thrifty-DAgger requires additional data collection and training for risk estimation, and Diff-DAgger achieves corresponding reduction factors for training, threshold setting, and inference times of 6.9, 24.9, and 8.8, respectively.
Figure 4: Average duration for training, threshold setting, inference, and total iteration for different methods, when training a visuomotor policy for Stacking. A single A100 40 GB GPU is used. The initial and final number of demonstrations $N_{i}$ and $N_{f}$, respectively, are 20 and 52, and training occurs after every $N_{d}$ = 4 interventions, with a total of 9 iterations.
@@ IV-E Discussions
False positives and false negatives reported in Table II have distinct implications: a false positive means the robot unnecessarily asks for expert demonstration, simply losing the benefits of interactive learning and resembling offline imitation learning. In contrast, a false negative, where the robot overestimates its ability and fails, is more dangerous, as it can cause the robot to deviate significantly from desired states over time, diminishing the value of expert intervention data. Both Ensemble-DAgger and Thrifty-DAgger incur a high rate of false negative in most tasks and we speculate that this is the reason they perform worse than behavior cloning in many tasks. On the other hand, Diff-DAgger yields far fewer false negatives than other methods across tasks, making it a more suitable choice for an interactive learning framework. This increased sensitivity to task failure is visualized in Fig 5, where we show robot query behavior from each DAgger method. For instance, Diff-DAgger proactively requests expert intervention once the robots start to lose grip of the charger, while the other methods fail to detect this issue—resulting in either dropping the object or getting stuck in one motion.
We also observe an interesting emergent behavior from policies trained on Diff-DAgger collected data: recovery behavior. As shown in Fig 6, after the policy fails at its initial attempt to stack the cube, it subsequently tries again and completes the task. This type of behavior is rarely seen when training the policy from offline behavior cloning, highlighting the effectiveness of Diff-DAgger.
@@
Figure 6: Emergent behavior of DAgger-trained policy for stacking. We observed an interesting new behavior with Diff-DAgger trained policies: the recovery behavior after initial failure. This explains the improvement in policy performance over BC-trained policies.
@@ V Conclusion & Future Work
In this work, we present Diff-DAgger, an efficient robot-gated DAgger method with diffusion policy. While Diff-DAgger shows improvements in task failure prediction, policy performance, and wall-clock time for learning over other methods, there are limitations: Whereas we dramatically reduce the wall-clock time for DAgger iteration, using an expressive policy class such as diffusion policy comes at the cost of increased training time in general. In addition, most of the choices we made to accelerate policy training is a trade-off between the policy performance, and we hope to address this in the future. Other future work includes learning from human experts for challenging real-world tasks. We hope that this work paves a way to open up more discussion for incorporating expressive policies into interactive learning frameworks in anticipation that computing will become faster over time.
@@ Acknowledgments
This work was funded and supported by Delta Electronics Inc.

### v4

@@ IV-C Task Performance
We evaluate the policy performance of methods for state-based and image-based observations. Diff-DAgger method consistently outperforms the baselines in task completion rate across tasks, assuming the same number of expert demonstrations is given. On average, Diff-DAgger achieves a 13.7% improvement over the second-best method in state-based tasks. In visuomotor tasks, including the real-world tasks, Diff-DAgger on average achieves a 25.2% improvement over the other methods.
@@ IV-D Wall-Clock Time for DAgger
We demonstrate the results on the visuomotor pushing task, where Diff-DAgger reduced training time by a factor of 5.1, threshold setting time by a factor of 48, and inference time by a factor of 3.8 compared to Ensemble-DAgger, as shown in Fig 2. The overall reduction in average policy update time is by a factor of 7.8, reducing the total time to just 0.54 hours. Thrifty-DAgger requires additional data collection and training for risk estimation, and Diff-DAgger achieves corresponding reduction factors for training, threshold setting, and inference times of 6.4, 53, and 3.9, respectively.
Figure 4: Average duration for training, threshold setting, inference, and total iteration for different methods, when training the final DAgger iteration of visuomotor policy for pushing task. A single A100 40 GB GPU is used for both training and inference.
@@ IV-E Discussions
False positives and false negatives reported in Table IV have distinct implications: a false positive means the robot unnecessarily asks for expert demonstration, simply losing the benefits of interactive learning and resembling offline imitation learning. In contrast, a false negative, where the robot overestimates its ability and fails, is more detrimental, as it can cause the robot to deviate significantly from the relevant states. Both Ensemble-DAgger and Thrifty-DAgger incur a high rate of false negative in most tasks, especially when learning from two pushing experts. This is likely because the action variance present in the training dataset is high due to diverging plans (clockwise paths and counter-clockwise paths) given the same state, resulting in the policies’ insensitivity toward true OOD data. And we speculate that this is the reason they perform worse than behavior cloning in many tasks. On the other hand, Diff-DAgger yields far fewer false negatives than other methods across tasks, making it a more suitable choice for an interactive learning framework with a complex data distribution. The sensitivity to task failure is visualized in Fig 5, where we show robot query behavior from each DAgger method. For instance, Diff-DAgger proactively requests expert intervention once the robots start to lose grip of the charger, while the other methods fail to detect this issue—resulting in either dropping the object or getting stuck in one motion.
It has been observed that the prediction type used for the diffusion model during training can have significant impact on both the convergence rate and the policy performance [29]. We find that sample prediction and V-objective prediction converges faster than epsilon and report the results using V-objective across the tasks as it works well considering both convergence rate and policy performance.
We also observe an interesting emergent behavior from policies trained on Diff-DAgger collected data: recovery behavior. As shown in Fig 6, after the policy fails at its initial attempt to stack the cube, it subsequently tries again and completes the task. This type of behavior is not typically observed unless carefully curated during data collection for offline behavior cloning.
@@
Figure 6: Emergent behavior of DAgger-trained policy for stacking. We observed an interesting new behavior with Diff-DAgger trained policies: the recovery behavior after initial failure. This explains the improvement in policy performance over policies trained using the offline approach.
@@ V Conclusion & Future Work
In this work, we present Diff-DAgger, an efficient robot-gated DAgger method with diffusion policy. While Diff-DAgger shows improvements in task failure prediction, policy performance, and wall-clock time for learning over other methods, there are limitations: Whereas we dramatically reduce the wall-clock time for DAgger iteration, using an expressive policy class such as diffusion policy comes at the cost of increased training time in general. In addition, most of the choices we made to accelerate policy training is a trade-off between the policy performance, and we hope to address this in the future. We hope that this work paves a way to open up more discussion for incorporating expressive policies into interactive learning frameworks in anticipation that computing will become faster over time.
@@ Acknowledgments
This work was funded and supported by Delta Electronics Inc. We acknowledge Research Computing at the University of Virginia for providing the computational resources and technical support.